# A compact neuromorphic system for ultra energy-efficient, on-device robot localization

### v1

@@
A compact neuromorphic system for ultra energy-efficient, on-device robot localization
Adam D Hines1∗, Michael Milford1, Tobias Fischer1
1QUT Centre for Robotics, Queensland University of Technology,
2 George St, Brisbane, QLD 4001, Australia
To whom correspondence should be addressed; E-mail: adam.hines@qut.edu.au
Neuromorphic computing offers a transformative pathway to overcome the computational and energy challenges faced in deploying robotic localization and navigation systems at the edge. Visual place recognition, a critical component for navigation, is often hampered by the high resource demands of conventional systems, making them unsuitable for small-scale robotic platforms which still require to perform complex, long-range tasks. Although neuromorphic approaches offer potential for greater efficiency, real-time edge deployment remains constrained by the complexity and limited scalability of bio-realistic networks. Here, we demonstrate a neuromorphic localization system that performs accurate place recognition in up to 8km of traversal using models as small as 180 KB with 44k parameters, while consuming less than 1% of the energy required by conventional methods. Our Locational Encoding with Neuromorphic Systems (LENS) integrates spiking neural networks, an event-based dynamic vision sensor, and a neuromorphic processor within a single SPECK™ chip, enabling real-time, energy-efficient localization on a hexapod robot. LENS represents the first fully neuromorphic localization system capable of large-scale, on-device deployment, setting a new benchmark for energy efficient robotic place recognition.
@@ Introduction
Robot localization is an essential component of any autonomous navigation system, enabling robots to not only determine their location but also understand and interact with their surroundings. A central challenge in robot localization is visual place recognition (VPR), which requires robots to identify and classify previously visited locations under varying conditions  (?, ?, ?). Conventional VPR strategies often rely on deep convolutional neural networks or transformer-based architectures to robustly extract features from the environment  (?, ?, ?, ?). For resource-constrained robotic platforms that nonetheless require the ability to keep track of where they are located over vast distances, using such conventional VPR methods is often not feasible due to their computational demands. An open challenge for real-world deployment of VPR is to find models that are computationally and energy efficient, and can be deployed at the edge.
@@
Neuromorphic computing, which takes inspiration by the brain, has emerged as a promising solution for addressing the energy and computational challenges associated with robot localization and navigation  (?, ?, ?, ?, ?). The human brain’s ability to learn and navigate complex environments efficiently  (?, ?, ?, ?) has inspired roboticists to create more compute efficient localization systems  (?, ?, ?, ?, ?). In particular, Spiking Neural Networks (SNNs) have been widely explored due to their ability to perform tasks through bio-realistic simulations of neuron activity  (?, ?). Whilst several SNN models have been proposed for robotic localization  (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?), they ultimately have not yet fully lived up to the promise of compute efficient deployment due to the complexity of continually modeling neurons in real-time  (?, ?, ?).
Figure 1: Neuromorphic place recognition system on a hexapod robot. Our system was deployed on a HiWonder Jetahexa hexapod robotic platform equipped with a custom 3D-printed mount housing a SynSense Speck2fDevKit (A). Event streams captured by the onboard DVS are processed into event frames by counting ON and OFF events during the hexapod’s traversal (B). The initial $128\times 128$ input is cropped to a $80\times 80$ region of interest which is then downsampled to $10\times 10$ by selecting central pixels via a 2D convolutional layer (see Materials and Methods: Event processing). (C) The spiking neural network, Locational Encoding with Neuromorphic Systems (LENS), learns reference event frames for on-chip deployment, enabling real-time localization using asynchronously collected events. (D) Sequence matching techniques enhance the similarity between reference and query inputs, improving overall precision. (E) Schematic of the fully neuromorphic visual place recognition pipeline.
@@
To advance the practicality of neuromorphic computing in VPR, the focus must shift toward developing models that can be effectively deployed on Size, Weight, and Power (SWaP) constrained robotic platforms. One approach to overcoming deployment barriers is to trade the intricate bio-realism of traditional SNNs for simpler, more efficient networks that still deliver robust performance. However, these small-scale networks remain limited in their ability to map large environments, which restricts the downstream use-cases  (?). To improve efficiency and scalability, SNNs full potential for robotic deployment can be realized by integrating them with neuromorphic hardware  (?, ?, ?, ?, ?, ?, ?, ?) such as Intel®’s Loihi 2, IBM’s TrueNorth, and Tianjic  (?, ?, ?), which draw inspiration from neuroscience for their design, enabling them to transmit and receive physical spikes from neurons within the processor cores  (?).
@@
Event cameras, including Dynamic Vision Sensors (DVS), offer further computational advantages when paired with neuromorphic processors as they only transmit information based on pixel-wise changes in light intensity exceeding a threshold, thereby reducing unnecessary data processing  (?). Fusing neuromorphic algorithms, sensors, and processors to perform robotic localization tasks therefore provides a promising avenue to overcome the computational limitations inherent in SNN-based VPR  (?, ?). While several neuromorphic localization systems have been developed and tested, they are limited by environment scale  (?), require complex multi-modal fusion of large network models  (?), or have restricted real-time capability  (?).
@@
Here, we introduce a novel neuromorphic pipeline for robotic localization that overcomes these limitations by enabling large-scale mapping and localization using a compact SNN with real-time performance. Our system, Locational Encoding with Neuromorphic Systems (LENS), is designed to deliver high accuracy and efficiency with a model size of less than 180 KB and 44K parameters, capable of processing up to 8 km of traversal data while consuming less than 1% of the energy required by conventional platforms. We train a temporally coded SNN designed to perform VPR  (?) and deploy our model on the SynSense SPECK™  (?), which combines a DVS and system-on-chip neuromorphic processor. By deploying on SPECK™, we forgo the need for any external sensory modalities (e.g. satellites) or compute resources as all localization is performed on-chip. By abstracting complex bio-realism in favor of enhanced performance, LENS represents a significant advancement in neuromorphic localization, as the first fully event-driven platform designed for VPR. We validate LENS on a HiWonder JetHexa hexapod robot (Fig. 1), demonstrating its effectiveness in both indoor and outdoor small-scale traversals ($\approx$25-40m, 50-80 places) and large-scale datasets ($\approx$8km,  600 places)  (?), showcasing its potential for online VPR.
@@ Results
Figure 2: Energy efficiency of Real-time VPR on a neuromorphic processor vs. von Neumann hardware. (A) Power consumption was recorded across 5 components of the SPECK™ neuromorphic processor: input/output (IO), memory (RAM), logic, digital components (VDDD), analog components (VDDA), and the total. Start indicates when the robot began to move until the navigation task has finished, indicated by the Stop. (B) The average power consumption rates for the SPECK™, Jetson Nano, and Intel® CPU were 3 mW, 536 mW, and 13,696 mW, respectively. In the context of our VPR task in (A), the SPECK™ consumed significantly less total energy compared to the Jetson Nano and Intel® CPU, consuming 327 mJ, 9865 mJ, and 377,858 mJ, respectively (C). The mW and mJ values are presented on a logarithmic scale for display purposes.
@@
Neuromorphic computing platforms offer a significant advantage in terms of power and energy consumption when compared to traditional von Neumann hardware, particularly in tasks like Visual Place Recognition (VPR) where energy efficiency is crucial for long-term operation on battery-powered robots. We measured the power consumed by various components of the SPECK™ neuromorphic processor (Fig. 2) while performing VPR, recorded in the range of milliwatts (Fig. 2A). Notably, when the robot is stationary, the DVS generates very few events which results in further reductions in power consumption. To quantify the energy savings of LENS deployed on the SPECK™ processor relative to von Neumann hardware, we simulated event streams for VPR using our LENS model on an Intel® i7-9700K CPU and a NVIDIA Jetson Nano, which runs the hexapod (see Materials and Methods: Off-chip simulation of event streams and Power measurements). We used simulated experiments to avoid additional energy costs associated with converting events from SPECK™ to event streams used for CPU processing. After subtracting baseline CPU power use (see Fig. S2 and S3), we found that the SPECK™ consumed an average of 2.7 mW, which is just 0.5% and 0.02% of the power required by the Jetson Nano and Intel® CPU, respectively, to perform the same operations (Fig. 2B).
@@
The overall energy consumption (the integral of power over time) of the SPECK™ was 327 mJ, with the Jetson using 9865 mJ and the Intel® CPU requiring 377,857 mJ, meaning overall our system required just 3.31% and 0.09% of the energy needed to perform the same task on von Neumann hardware (Fig. 2C).
@@
Figure 3: Precision of a real-time neuromorphic robotic localization platform. We deployed our robotic platform in two distinct areas, indoor (A) and outdoor (F), to compare how our system compares in different environments. Mapping and localization is performed through a predetermined route in each environment using the hexapod, with slight differences in viewpoint between the two datasets (B, G). To analyze the precision and recall of our model, we generated sequenced matched similarity matrices of the output spikes from all output neurons in our model while localizing (C, H). Our compact system performed comparably to the sum of absolute differences (SAD) method for Recall@N (D, I) and Precision-Recall (E, J), with LENS being better suited for indoor routes than outdoor.
@@
To evaluate our system’s real-time performance, we performed VPR on our hexapod in both indoor and outdoor environments (Fig. 3). The model architecture was designed with $100\times 200\times 75$ for the input, feature, and output neurons, respectively. As illustrated in Figs. 3A & F, our generated event frames correspond to the environment being traversed, and particularly detect edges in the scene. The robot dealt with viewpoint and timing differences between the mapping and localization phases, as the route paths differed due to imperfect teleoperation (Figs. 3B & G). We used spikes from the model’s output layer to generate similarity matrices (Figs. 3C & H, see Materials and Methods: Online place matching). The similarity matrix was used to compute precision and recall metrics, which evaluates how well the robot is correctly identifying its location as it moves through the environment (Fig. 3C & H). We also employed the Sum-of-Absolute-Differences (SAD) method  (?, ?) as a baseline comparison method, calculating pixel-wise similarity between mapping (training) and localization (test) data.
For the indoor task, LENS achieved a Recall@1 of 0.36, comparable to SAD’s performance at 0.35 (Fig. 3D). The precision-recall curve (PR-curve) indicated that LENS achieved higher precision at lower recall values compared to SAD (Fig. 3E), suggesting that LENS is highly accurate in its identification of places, even if it recognizes fewer places overall. In the outdoor task, however, SAD outperformed LENS with a Recall@1 of 0.17 and 0.15, respectively (Fig. 3I). SAD also achieved a higher precision at lower recall, shown in the PR-curve (Fig. 3J).
@@
The reduced performance in outdoor environments is likely due to the higher dynamic range and higher incidence of non-static objects. As indicated by the similarity matrix (Fig. 3H), LENS showed a higher amount of activity across multiple output neurons. Fig. S1 suggests that increased contrastive differences in lighting during the navigation task introduced noise, distributing spikes across more neurons.
@@
Overall, our compact spiking neural network model (130-150 KB in size) demonstrated comparable performance for VPR tasks in multiple environments, with key advantages of real-time deployability whilst being highly power and energy efficient. By contrast, the SAD method requires 388 KB to store the images, $\approx$2.6 times the storage space needed for our model. For scalable applications, our system offers clear benefits in terms of model size, making it feasible for deployment on resource-constrained compute platforms.
Large-scale place recognition with compact models
Figure 4: Large scale place recognition using a compact neuromorphic ecosystem. (A) Coordinate map of the Brisbane Event VPR dataset route  (?) which represents an $\approx$8km traversal and 641 unique place locations around the suburbs of Brisbane, Australia. (B) Event frames are generated in the same way as on robot experiments (Fig. 3). Pixels were selected randomly rather than applying a 2D convolutional filter due to a larger event camera resolution used in  (?). We selected pixels to generate event frames that are $7\times 7$ in size. Our model architecture was modified to $49\times 63\times 641$ to allow for on-chip inferencing. (C) Similarity matrices for both SAD and LENS, showing how the similarity for both methods aligns closely with the ground truth (GT). (D) Recall@N and (E) Precision-Recall curves for our method when compared to SAD. SAD performed better at Recall@1 with 0.63 compared to ours at 0.36, however we achieved better precision at lower recall indicating our system is more confident at the places it selected. The energy efficiency and compact model size of just 753 neurons with 44K parameters highlights the benefit of our system.
To assess our system’s suitability for other real-world scenarios, we evaluated its performance in large scale environments using the Brisbane Event VPR dataset containing event streams from an $\approx$8km route in Brisbane, Australia  (?) (Fig. 4). This dataset circumvents limitations due to the hexapod’s low movement speed (see Materials and Methods: Robotic deployment). We generated 641 event frames of unique places by sampling over 1 second intervals from the dataset (Fig. 4B). To fit the model onto the SPECK™ chip, we adjusted the network architecture to 49 input neurons, 63 feature layer neurons, and 641 output neurons. This was the largest and highest performing model architecture we could use to run our model on-chip, balancing the number of input pixels and feature representation (Fig. S4). We used the “sunset2” traverse for mapping (reference dataset), and the “sunset1” traverse for localization (query dataset).
Both LENS and SAD generated similarity matrices that closely aligned with the ground truth (Fig. 4C). Although SAD outperformed LENS for Recall@N, achieving a Recall@1 of 0.63 compared to LENS’s 0.36 (Fig. 4D), LENS demonstrated better precision at lower recall, indicating higher confidence in its place recognition (Fig. 4E). However, we note that a spiking neural network achieving this degree of accuracy with an architecture of just 753 neurons across 3 layers shows promise for large scale place recognition applications. We required the same amount of storage space capacity to store our model and the down sampled images for SAD, which was $\approx$179 KB. If memory constraints on SPECK™ had not been a limiting factor, larger LENS networks could perform as well or better than SAD (Fig. S4). For example, a model architecture of $49\times 98\times 641$ input, feature, and output neurons respectively measures at just 278 KB and performs comparably to SAD (Fig. S4).
@@
Interestingly, both SAD and LENS performed better on the Brisbane Event VPR dataset than during the on-robot VPR task (Fig. 3). The Brisbane Event Dataset utilized a $346\times 240$ sensor  (?), compared to the $128\times 128$ sensor  (?) on our robot. Additionally, the more stable platform (i.e., a car) used for dataset collection could reduce errors from viewpoint variance or swaying movement observed with the hexapod (see Movie 1). These results indicate that our proposed model can be effectively deployed in various environments and on different robotic platforms for a wide range of localization and navigation tasks.

### v2

# A compact neuromorphic system for ultra-energy-efficient, on-device robot localization

@@
Abstract
Neuromorphic computing offers a transformative pathway to overcome the computational and energy challenges faced in deploying robotic localization and navigation systems at the edge. Visual place recognition, a critical component for navigation, is often hampered by the high resource demands of conventional systems, making them unsuitable for small-scale robotic platforms which still require accurate long-endurance localization. Although neuromorphic approaches offer potential for greater efficiency, real-time edge deployment remains constrained by the complexity of bio-realistic networks. In order to overcome this challenge, fusion of hardware and algorithms is critical to employ this specialized computing paradigm. Here, we demonstrate a neuromorphic localization system that performs competitive place recognition in up to 8 kilometers of traversal using models as small as 180 kilobytes with 44,000 parameters, while consuming less than 8% of the energy required by conventional methods. Our Locational Encoding with Neuromorphic Systems (LENS) integrates spiking neural networks, an event-based dynamic vision sensor, and a neuromorphic processor within a single SynSense Speck™ chip, enabling real-time, energy-efficient localization on a hexapod robot. When compared to a benchmark place recognition method, Sum-of-Absolute-Differences (SAD), LENS performs comparably in overall precision. LENS represents an accurate fully neuromorphic localization system capable of large-scale, on-device deployment for energy efficient robotic place recognition. Neuromorphic computing enables resource-constrained robots to perform energy efficient, accurate localization.
@@ Introduction
Robot localization is a critical component of many autonomous navigation systems, enabling robots to determine their location whilst supporting the ability to understand and interact with their surroundings. A central challenge in robot localization is visual place recognition (VPR), which requires robots to identify and classify previously visited locations under varying conditions  (?, ?, ?). Conventional VPR strategies often rely on deep convolutional neural networks or transformer-based architectures to robustly extract features from the environment  (?, ?, ?, ?). For resource-constrained robotic platforms that nonetheless require the ability to keep track of where they are located over vast distances, using such conventional VPR methods is often not feasible due to their computational demands. An open challenge for real-world deployment of VPR is to find models that are computationally and energy efficient, and can be deployed at the edge.
@@
Neuromorphic computing, which takes inspiration by the brain, has emerged as a promising solution for addressing the energy and computational challenges associated with robot localization and navigation  (?, ?, ?, ?, ?). The human brain’s ability to learn and navigate complex environments efficiently  (?, ?, ?, ?) has inspired roboticists to create more computationally efficient localization systems  (?, ?, ?, ?, ?). In particular, Spiking Neural Networks (SNNs) have been widely explored due to their ability to perform tasks through bio-realistic simulations of neuron activity  (?, ?). Although several SNN models have been proposed for robotic localization  (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?), they ultimately have not yet fully lived up to the promise of computationally efficient deployment due to the complexity of continually modeling neurons in real-time  (?, ?, ?).
@@
To advance the practicality of neuromorphic computing in VPR, the focus must shift toward developing models that can be effectively deployed on Size, Weight, and Power (SWaP) constrained robotic platforms. One approach to overcoming deployment barriers is to trade the intricate bio-realism of traditional SNNs for simpler, more efficient networks that still deliver robust performance  (?). However, small-scale systems can often be limited in their ability to map large environments, which restricts the downstream use-cases  (?). To improve efficiency, SNNs full potential for robotic deployment can be realized by integrating them with neuromorphic hardware  (?, ?, ?, ?, ?, ?, ?, ?) such as Intel®’s Loihi 2, TrueNorth, and Tianjic  (?, ?, ?), which draw inspiration from neuroscience for their design, enabling them to transmit and receive physical spikes from neurons within the processor cores  (?). Neuromorphic computing has been successfully utilized in tasks other than localization, such as control systems for drone flight  (?), route following  (?), autonomous driving controllers  (?), and motor controllers  (?). In contrast, non-spiking algorithms that are brain-inspired are widely used in localization tasks but have limited ability to be deployed on neuromorphic hardware. Although several neuromorphic localization systems have been developed and tested, they are limited by environment scale  (?), require complex multi-modal fusion of large network models  (?), or have restricted real-time capability  (?).
@@
Event cameras, including Dynamic Vision Sensors (DVS), offer further computational advantages when paired with neuromorphic processors as they only transmit information based on pixel-wise changes in light intensity exceeding a threshold, thereby reducing unnecessary data processing  (?). In the context of SNNs, spiking activity can easily be triggered and propagated based on incoming input event streams and are well suited to process them. Fusing neuromorphic algorithms, sensors, and processors to perform robotic localization tasks therefore provides a promising avenue to overcome the computational limitations inherent in SNN-based VPR  (?, ?).
@@
Here, we introduce a neuromorphic pipeline for robotic localization that is compact, real-time capable, and able to map environments up to 8 km in length. Our system, Locational Encoding with Neuromorphic Systems (LENS), is designed to deliver high accuracy and efficiency with a model size of less than 180 KB and 44 K parameters, capable of processing up to 8 km of traversal data while consuming less than 8% of the energy required by conventional platforms. We train a SNN using a temporal time to first spike encoding scheme designed to perform VPR  (?) and deploy our model on the SynSense Speck™  (?), which combines a DVS and system-on-chip neuromorphic processor. The LENS neuromorphic algorithm was developed in tandem with neuromorphic hardware deployment to deliver ultra energy-efficient place recognition for robotic localization. By deploying on Speck™, we forgo the need for any external sensory modalities or computational resources as all localization is performed on-chip. By abstracting complex bio-realism in favor of enhanced performance, LENS represents a notable advancement in neuromorphic localization, as a fully event-driven platform designed for VPR. We validate LENS on a HiWonder JetHexa hexapod robot (Fig. 1), demonstrating its effectiveness in both indoor and outdoor small-scale traversals ($\approx$25-40 m, 50-80 places) and large-scale datasets ($\approx$8 km,  600 places)  (?), showcasing its potential for online VPR.
Figure 1: Neuromorphic place recognition system on a hexapod robot. (A) Our system was deployed on a HiWonder Jetahexa hexapod robotic platform equipped with a custom 3D-printed mount housing a SynSense Speck2fDevKit. Event streams captured by the onboard DVS are processed into event frames by counting ON and OFF events during the hexapod’s traversal (B). The initial $128\times 128$ input is cropped to a $80\times 80$ region of interest which is then downsampled to $10\times 10$ by selecting central pixels via a 2D convolutional layer (see Materials and Methods: Event processing). (C) The spiking neural network, Locational Encoding with Neuromorphic Systems (LENS), learns reference event frames for on-chip deployment, enabling real-time localization using asynchronously collected events. (D) Sequence matching techniques enhance the similarity between reference and query inputs, improving overall precision. (E) Schematic of the fully neuromorphic visual place recognition pipeline.
@@ Results
Locational encoding using neuromorphic systems
An overview of our Locational Encoding using Neuromorphic Systems (LENS) network is presented in Figure 2. We used event streams from the Speck™ DAVIS128 sensor in two different ways for the mapping and localization phases of the robot navigation. A place was represented as one second of movement, translating to a distance of $\approx$0.15 m (Fig. 2A). For the mapping phase, the number of events were counted per pixel and then flattened into a 2D-temporal representation image of the place (Fig. 2B). Using a 2D-convolutional kernel with the center weight set to 1 and all other weights set to 0 (see Materials and Methods: Model training), we selected the center pixel from the 2D image to downsample to $10\times 10$ pixels to fit in our LENS system (Fig. 2B). Following Hines et al.  (?), we trained our model by first normalizing pixel intensities between [0, 1] to use a temporally coded spike scheme such that higher pixel intensities result in earlier spiking with lower intensities activating later (Fig. 2C, see Materials and Methods: Model training). Connection weights were trained via spike-timing dependent plasticity learning, with enhanced connectivity when post-synaptic activity occurred after pre-synaptic spikes and diminished connectivity if post-synaptic neurons spike beforehand (Fig. 2C, see Materials and Methods: Model training). For our three layer architecture, we used unsupervised learning between the input and feature layers and a supervised delta learning rule between the feature and output layer (see Materials and Methods: Model training).
Figure 2: Event stream processing using LENS. (A) Event streams are collected from the Spec2fDevKit board from an $80\times 80$ region of interest using a DAVIS128 sensor as our hexapod robot moves through the environment. One place represents one second or $\approx$0.15 m of movement. The event stream is then used in two unique ways during mapping and localization phases. (B) During mapping, events over the one second time period are counted and flattened into a 2D-temporal representation of the place. We use a convolutional selection kernel to pick the center pixel from the 2D image (see Materials and Methods: Model training), resulting in a $10\times 10$ down sampled image. (C) Down sampled images are normalized in the range of $[0,1]$ with pixel intensity determining spike timing in our LENS training regime, following  (?). Using spike-timing dependent plasticity (STDP) across the network layers, we strengthen or weaken both excitatory and inhibitory connections if the spike occurs after or before pre-synaptic activity, respectively. (D) During localization, we forgo the temporal representation of places and use the raw event streams to continuously activate and bias the LENS network on-chip. (E) We apply the same convolutional pixel selection as the first layer of our neural network, deployed on the Speck™ directly. Over the time collection period of one second, we accumulate output spikes from our model and make a place prediction based on the neuron with the highest spiking activity.
During the localization phase, we used the raw event streams collected from the Speck™ directly into our model instead of 2D-temporal representations of places (Fig. 2D, see Materials and Methods: Deployment time). The first layer of our LENS system was the convolutional pixel selection filter applied earlier to reduce the input dimensionality to $10\times 10$ pixels (Fig. 2E, see Materials and Methods: Deployment time). We performed place recognition with asynchronous events over one second collection windows and picked the location with the highest spiking activity for the neuron in the output layer (Fig. 2E, see Materials and Methods: Online place matching).
@@
Neuromorphic computing platforms offer an advantage in terms of power and energy consumption when compared to traditional von Neumann hardware, particularly in tasks like Visual Place Recognition (VPR) where energy efficiency is crucial for long-term operation on battery-powered robots. We measured the power consumed by various components of the Speck™ neuromorphic processor while performing VPR (Fig. 3A). Notably, when the robot was stationary, the DVS generated very few events which resulted in further reductions in power consumption. To quantify the energy savings of LENS deployed on the Speck™ processor relative to von Neumann hardware, we performed VPR using the Sum-of-absolute-differences (SAD) place matching method  (?) on an Intel® i7-9700K CPU and a NVIDIA Jetson Nano, which ran the hexapod (see Materials and Methods: Power measurements). After subtracting baseline CPU power use, we found that the Speck™ consumed an average of 2.7 mW, which is just 8.8% and 0.5% of the power required by the Jetson Nano and Intel® CPU to run SAD, respectively (Fig. 3B).
Figure 3: Energy efficiency of Real-time VPR on a neuromorphic processor vs. von Neumann hardware. (A) Power consumption was recorded across 5 components of the Speck™ neuromorphic processor: input/output (IO), memory (RAM), logic, digital components (VDDD), analog components (VDDA), and the total. Start indicates when the robot began to move until the navigation task has finished, indicated by the Stop. (B). The Speck™ processor consumed substantially less total energy compared to the Jetson Nano and Intel® CPU (C). The mW and mJ values are presented on a logarithmic scale for display purposes.
@@
The overall energy consumption (the integral of power over time) of the Speck™ was 327 mJ, with the Jetson using 2,968 mJ and the Intel® CPU requiring 61,427 mJ, meaning overall our system required just 8.9% and 0.5% of the energy needed to perform the same task on von Neumann hardware (Fig. 3C). Fewer pixel representations of places still resulted in energy consumption far greater than what was required for LENS on Speck™ (see Table S2). We also observed a prominent reduction in power and energy consumption deploying to Speck™ when running our LENS system on von Neumann hardware (Fig. S1).
Large-scale place recognition with compact models
To assess our system’s suitability for real-world scenarios, we evaluated its performance in large scale environments using the Brisbane Event VPR dataset containing event streams from an $\approx$8 km route in Brisbane, Australia  (?) (Fig. 4). This dataset was representative of a stable, four wheeled robotic platform that our model could be deployed on. We generated 641 event frames of unique places by sampling over one second intervals from the dataset (Fig. 4B). To fit the model onto the Speck™ chip, we used a network architecture of 49 input neurons, 63 feature layer neurons, and 641 output neurons (Fig. 4B). This was the largest and highest performing model architecture we could use to run our model on-chip, balancing the number of input pixels and feature representation (Fig. S6). We used the “sunset2” traverse for mapping (reference dataset), and the “sunset1” traverse for localization (query dataset). We also employed the Sum-of-absolute-difference (SAD method)  (?, ?) as a baseline comparison method, calculating pixel-wise similarity between mapping (training) and localization (test) data.
Figure 4: Large scale place recognition using a compact neuromorphic ecosystem. (A) Coordinate map of the Brisbane Event VPR dataset route  (?) which represents an $\approx$8 km traversal and 641 unique place locations around the suburbs of Brisbane, Australia. (B) Event frames are generated in the same way as on robot experiments (Fig. 5). Pixels were selected randomly rather than by applying a 2D convolutional filter due to a larger event camera resolution used in  (?). We selected pixels to generate event frames that are $7\times 7$ in size. Our model architecture was modified to $49\times 63\times 641$ to allow for on-chip localization. (C) Similarity matrices for both SAD and LENS, showing how the similarity for both methods aligns closely with the ground truth (GT). (D) Recall@N and (E) Precision-Recall curves for our method when compared to SAD. LENS performed best overall compared to SAD, and also achieved better precision at lower recall indicating our system is more confident at the places it selected. The energy efficiency and compact model size of just 753 neurons with 44K parameters highlights the benefit of our system.
We used spikes from the model’s output layer and the Euclidean distance calculated in SAD to generate similarity matrices and used sequence matching which preserves the distances of sequences matches while nullifying cross-distances, improving localization precision during deployment time (Fig. 4C, see Materials and Methods: Online place matching). The similarity matrix was used to compute Recall@N (the proportion of true place matches retrieved within the top N predicted matches) and precision and recall metrics, which evaluated how well the robot is correctly identifying its location as it moves through the environment. Although other methods for event based place recognition exist, they require the reconstruction of images from event streams for deep feature extraction  (?, ?). As such, we elected to not compare our system to these methods since they require higher complexity and computational time than what is feasible to run on our robotic platform.
Both LENS and SAD generated similarity matrices that closely aligned with the ground truth (Fig. 4C). LENS performed best overall for Recall@N, achieving a Recall@1 of 0.88 compared to SAD’s 0.81 (Fig. 4D). LENS also demonstrated better precision at lower recall, indicating higher confidence in its place recognition (Fig. 4E). The non-monotonic behavior in some precision-recall curves, where precision temporarily improves with increasing recall, occured when the rate of true positive matches increased faster than false positives at certain threshold ranges. A spiking neural network to achieve this degree of accuracy with an architecture of just 753 neurons across three layers shows promise for large scale place recognition applications. We required the same amount of storage space capacity to store our model and the down sampled images for SAD, which was $\approx$179 KB.
@@
To evaluate our system’s real-time performance at the edge, we performed VPR on our hexapod in both indoor and outdoor environments (Fig. 5). The model architecture was designed with $100\times 200\times 75$ for the input, feature, and output neurons, respectively. As illustrated in Figs. 5A & F, the generated event frames corresponded to the environment being traversed, particularly by detecting edges in the scene. The robot dealt with viewpoint and timing differences between the mapping and localization phases, as the route paths differed due to imperfect teleoperation (Figs. 5B & G). In comparison to the Brisbane Event Dataset, the hexapod represented a more unstable and challenging platform to perform place recognition.
Figure 5: Precision of a real-time neuromorphic robotic localization platform. We deployed our system on a Hexapod robotic platform for on-device localization. (A) First, we tested an indoor environment by mapping and localizing along a predetermined route (B). As the Hexapod was teleoperated, there are slight differences in viewpoint between the mapping and localization phases. (C) To analyze the precision and recall of our indoor environment, we generated sequentially matched similarity matrices of the output spikes from all output neurons in our model while localizing. (D) Our compact system generally performed better than the sum of absolute differences (SAD) method for Recall@N, measuring accuracy in the top N matches, and Precision-Recall (E). We then deployed our Hexapod in an outdoor environment (F, G). The similarity matrix for the outdoor environment had more aliased incorrect matches than the indoor one (H), most likely due to the presence of multiple turns in the route. (I, J) In the outdoor task, LENS performed comparably to the SAD method in both the Recall@N and Precision-Recall evaluation.
@@
For the indoor task, LENS achieved a Recall@1 of 0.64, outperforming SAD’s performance at 0.36 (Fig. 5C & D). The precision-recall curve (PR-curve) indicated that LENS achieved higher precision at lower recall values compared to SAD (Fig. 5E), suggesting that LENS was highly accurate in its identification of places. In the outdoor task, LENS performed comparably to SAD with a Recall@1 of 0.33 and 0.29, respectively (Fig. 5H & I). LENS also achieved a higher precision at lower recall, shown in the PR-curve (Fig. 5J). The reduced performance in outdoor environments was likely due to the increased amount of turns in the trajectory, which caused widespread neuronal activation across the field of view. As indicated by the similarity matrix (Fig. 5H), LENS showed a higher amount of activity across multiple output neurons. Fig. S8 suggests that increased turning in the navigation route introduced noise, distributing spikes across more neurons.
@@
Overall, our compact spiking neural network model (130-150 KB in size) demonstrated comparable performance for VPR tasks in multiple environments, with key advantages of real-time deployability whilst being highly power and energy efficient. By contrast, the SAD method required 388 KB to store the images, $\approx$2.6 times the storage space needed for our model. Our system offered substantial benefits in terms of model size, making it feasible for deployment on computationally resource-constrained platforms.
@@
Interestingly, both SAD and LENS performed better on the Brisbane Event VPR dataset than during the on-robot VPR task (Fig. 5). The Brisbane Event Dataset utilized a $346\times 240$ sensor  (?), compared to the $128\times 128$ sensor  (?) on our robot. Additionally, the more stable platform (i.e., a car) used for dataset collection could reduce errors from viewpoint variance or swaying movement observed with the hexapod (see Supplementary Movie 1). These results indicate that our proposed model can be effectively deployed for both short and long-range place recognition tasks for robotic localization.